Good day,

A fuel was used to heat water in a calorimetry experiment. Propane was burnt causing a temperature change of 2.75 K in 700 g of water (specific heat capacity = 4.18 J/g·K).


Calculate the heat change of the water.


Heat change (J) = mass (g) × specific heat capacity (J/g·K) × temperature change (K)


Give your answer to three significant figures.    Thanks


1
Expert's answer
2020-08-20T04:01:45-0400

As per the question

"\triangle T=2.75 K"

Specific heat "(C)=4.18 J\/g.K"

Mass of water "(m)= 2.75\ g"

Heat change in water"= mC\triangle T= 2.75\times 4.18\times 700=8046.5 J"

"=8.05\times 10^3J" (in 3 significant figures)